OpenStudy (anonymous):

A shirt originally cost $40.98, but it is on sale for $36.88. What is the percentage decrease of the price of the shirt? If necessary, round to the nearest percent. 90% 10% 11% 8%

OpenStudy (phi):

first figure out change in price divided by original price

OpenStudy (anonymous):

36.88/40.98= 0.89 (89%) Then subtace it from 100 to find decrease 100-89= 11 11% 36.88x1.11= 40.9368 (round to 40.98)

OpenStudy (anonymous):

Jane moved from a house with 61 square feet of closet space to an apartment with 41.48 square feet of closet space. What is the percentage decrease of Jane's closet space? 47% 68% 32% 26%

OpenStudy (anonymous):

41.48/61=0.68 (68%) 100-68=32 32% decrease :)

OpenStudy (phi):

@madicayer These percent change problems can be tricky. The definition is (original - new)/original changed to a percent see http://www.wikihow.com/Calculate-Percent-Change

OpenStudy (anonymous):

i am really stuck on this one In his free time, Gary spends 9 hours per week on the Internet and 7 hours per week playing video games. If Gary has five hours of free time per day, approximately what percent of his free time is spent on the Internet and playing video games? 46% 3% 13% 32%